OpenStudy (anonymous):

Find the area of the larger segment whose chord is 8" long in a circle with an 8" radius. (Hint: A chord divides a circle into two segments. In problem 1, you found the area of the smaller segment.)

OpenStudy (anonymous):

Area of a circle is pi times radius squared. so 8 squared is 64 times 3.14 is 200.96. 200.96 would be the approx. answer

OpenStudy (anonymous):

The answer has to be in this type of way: A=(XXX/X PI + XX sqrt X) in sq

OpenStudy (cwrw238):



OpenStudy (cwrw238):

you can find the area of smaller segment by finding the area of the smaller sector - area triangle then subtract this from area of the circle

OpenStudy (cwrw238):

the triangle is equilateral so the angle at the centre is pi/3 radians

OpenStudy (cwrw238):

area of sector = 0.5 * r^2 * (pi/3) area of triangle = 0.5 * 8*8* sin (pi/3)

OpenStudy (cwrw238):

oh - i'm talking to myself again!